OpenStudy (lgbasallote):

\[(x \text D + 2)(x\text D - 1)\]

OpenStudy (lgbasallote):

what's its difference from \[(x\text D - 1)(x\text D + 2)\] then?

OpenStudy (lgbasallote):

something smells fishy :/

OpenStudy (lgbasallote):

ohhh i see it

OpenStudy (lgbasallote):

there's gonna be xD2 instead of 2xD

OpenStudy (anonymous):

no it is wrong..

OpenStudy (lgbasallote):

wait what o.O

OpenStudy (anonymous):

Wait..

OpenStudy (lgbasallote):

it shouldnt be the same o.O

OpenStudy (lgbasallote):

there are x's here

OpenStudy (anonymous):

No.. they are not same..

OpenStudy (lgbasallote):

this is what i think \[(x\text D + 2)(x\text D - 1) \implies x^2 \text D^2 + x\text D(-1) + 2x \text D - 2\] i think you made a mistake o.O

OpenStudy (anonymous):

\[(xD−1)(xD+2) \implies x^2D^2 + xD(2) -xD - 2 \implies x^2D^2 -xD - 2\]

OpenStudy (anonymous):

It is the second one..

OpenStudy (lgbasallote):

so then that gives me \[x^2 \frac{d^2y}{dx^2} - x + 2x \frac{dy}{dx} - 2\] right?

OpenStudy (lgbasallote):

yeah i was saying i think you made a mistake in the original question

OpenStudy (anonymous):

\[\large = x^2D^2 +xD(-1) + 2xD - 2 \implies x^2D^2 + 2xD - 2\] This is the original one..

OpenStudy (lgbasallote):

what happened to -x?

OpenStudy (anonymous):

D(constant) = 0

OpenStudy (lgbasallote):

oh yeah...how can i forget

OpenStudy (anonymous):

So the two DEs are different.. I don't know the name of person but he told you that they never commutate.. Remember this for lifetime..